Nancy left downtown Paris, and three hours later, Sam left going 67 mph faster to catch up. After another two hours, Sam caught up. Find Nancy's average speed.
Found 2 solutions by Boreal, stanbon:
Answer by Boreal(15235)

Nancy's speed is x mph
Sam's is x+67 mph
Nancy traveled a total of 5 hours
Sam traveled 2 hours
distance is time* speed; distances are equal, since Nancy was caught.
5x=2(x+67)
5x=2x+134
3x=134
x=44 2/3 mph for Nancy; x+67=111 2/3 mph for Sam
Check
Sam left after Nancy had traveled 134 miles (44 2/3 times 3)
He was traveling at 67 mph faster than she.
Therefore, he caught her in 2 hours.

Answer by stanbon(75887)

Nancy left downtown Paris, and three hours later, Sam left going 67 mph faster to catch up. After another two hours, Sam caught up. Find Nancy's average speed.
=================
Nancy DATA:
distance = x miles ; time = 5 hrs ; rate = x/5 mph
------------
Sam DATA:
distance = x miles ; time = 2 hrs ; rate = x/2 mph
------------
Equation:
x/2 = 67 mph
x = 134 miles
-------
Ans: x/5 = 134/5 = 26.8 mph (Nancy's speed)
---------------
Cheers,
Stan H.
-------------